220 v dc shunt motor has Ra= 0.25 ohms and Rf=120 ohms .As7p44yy4t8GAAAAAElFTkSuQmCC At no-load the motor draws a current It=6A.x7HcTct+NNqclIKPKHOdO7tmvLfowOa6G2zvWeq4 And runs at 1500 rpm. At full-load the motor draws 62 A. the armature reaction at full load reduces the flux by 10%.

Determine the rotational losses , motor speed at full-load, the shaft torque and efficiency.
Answer #1

The page number is given on the top corners of each page:

page Solution Given : V = 220v, Ra=0.252, Rf=1202 - Load It = 6A , tt Speed (No) = 15000.p.m. - 6 * At Full load It = 62 andpage ② page : Rotational losses (Pr ) = Eola (At no load) = (2.18.95) x 4.17 walt I P = 913.02 Walt Ansi since Ex 0 N ; N= sppage power output = Elace Po Py = Rotational losses ... Poutput = (204.95)*(60-17) - 913.02. = 11418.82 Watt Pout = 11.42 kW]